# Uncovering Vulnerabilities of LLM-Assisted Cyber Threat Intelligence

### v2

@@ Abstract
Large Language Models (LLMs) are intensively used to assist security analysts in counteracting the rapid exploitation of cyber threats, wherein LLMs offer cyber threat intelligence (CTI) to support vulnerability assessment and incident response. While recent work has shown that LLMs can support a wide range of CTI tasks such as threat analysis, vulnerability detection, and intrusion defense, significant performance gaps persist in practical deployments. In this paper, we investigate the intrinsic vulnerabilities of LLMs in CTI, focusing on challenges that arise from the nature of the threat landscape itself rather than the model architecture. Using large-scale evaluations across multiple CTI benchmarks and real-world threat reports, we introduce a novel categorization methodology that integrates stratification, autoregressive refinement, and human-in-the-loop supervision to reliably analyze failure instances. Through extensive experiments and human inspections, we reveal three fundamental vulnerabilities: spurious correlations, contradictory knowledge, and constrained generalization, that limit LLMs in effectively supporting CTI. Subsequently, we provide actionable insights for designing more robust LLM-powered CTI systems to facilitate future research.
@@ 1 Introduction
We are living in an era of rapid digital transformation, where technological advancements are tightly associated with the growing prevalence of cyber threats. In recent years, the cyber threat landscape has shifted dramatically, with reported Common Vulnerabilities and Exposures (CVEs) increasing by an average of 25% annually (Intelligence, 2023). In 2024 alone, more than 40,000 vulnerabilities were reported (Corporation, 2025). This surge can be attributed to the rising complexity of IT systems (Qualys, 2024), the widespread adoption of open-source software (Dam & Neumaier, 2023; Reading, 2024), and the accelerating pace of modern development cycles (Security, 2024). Together, these dynamics expand the attack surface while making sole reliance on human analysts for vulnerability assessment and remediation increasingly infeasible.
Figure 1: (Left) Failure ratios of cybersecurity agents. (Right) Examples of vulnerabilities.
@@
Large Language Models (LLMs) have recently demonstrated strong performance in a broad range of cyber threat intelligence (CTI) tasks. By adapting models through instruction fine-tuning or prompt-based automation, researchers have applied LLMs to support threat analysis and decision-making (Zhang et al., 2023), code vulnerability detection (Du et al., 2024), and defense against network intrusions (Lavi et al., 2024). Despite these advances, substantial performance gaps remain in their reported evaluations (Deng et al., 2024; Clairoux-Trepanier et al., 2024; Ji et al., 2024; Alam et al., 2024; Liu et al., 2025), suggesting that such limitations cannot be fully addressed through typical model adaptation or prompt-based automation. This raises a fundamental research question: What intrinsic vulnerabilities constrain the effectiveness of LLMs in supporting CTI tasks?
Although some vulnerabilities stem from general challenges in LLM architecture and training methodologies (which affect not only CTI but also broader domains (Aguilera-Martínez & Berzal, 2025)), this work focuses on the nature of the threat landscape itself, which may introduce unique vulnerabilities for LLMs in effectively supporting CTI. Specifically, CTI requires reasoning under intertwined, crowdsourced, and imbalanced evidence, which are conditions that differ substantially from standard NLP benchmarks. To investigate this, this work makes the following contributions:
First, we conduct large-scale, systematic studies of LLM vulnerabilities in CTI, spanning a full-fledged CTI lifecycle from threat analysis to incident responses. Our evaluation datasets drawn from multiple CTI benchmarks (CTIBench (Alam et al., 2024), SevenLLM-Bench (Ji et al., 2024), SWE-Bench (Jimenez et al., 2023), and CyberTeam (Liu et al., 2025)) as well as real-world threat reports from CTI databases (CVE Program, 2024; National Institute of Standards and Technology (2024), NIST; Cybersecurity and Infrastructure Security Agency (2024), CISA) to ensure comprehensiveness.
Next, we develop a novel categorization methodology to investigate failure instances. One major challenge in studying extensive threat instances is reliably scaling the analysis to categorize diverse failure cases. Standard “LLM-as-judge” approaches are unreliable for automatic vulnerability type finding and classification, as models tend to rationalize their own reasoning or fail to critically evaluate contradictions (Yamauchi et al., 2025; Guerdan et al., 2025). To address this, we propose an autoregressive, human-in-the-loop framework that categorizes failure instances efficiently and with high reliability (§3). This methodology allows us to extract consistent insights from large-scale evaluations while minimizing annotation workload.
@@
Figure 2: Summarization of the vulnerability types of LLMs in various CTI stages. Ratios are calculated over the entire dataset (§2.2). Vulnerabilities may overlap within a single threat instance.
Lastly, through large-scale experiments on industry-leading LLMs and LLM-powered cybersecurity agents, complemented by detailed case studies, we identify three dominant categories of vulnerabilities that limit LLMs in handling the threat landscape, as illustrated in Figure 1:
(1) Spurious correlations, wherein LLMs over-attribute based on superficial or co-occurring features (e.g., mistaking commodity tools such as Mimikatz (Delpy, 2011) as actor-specific evidence).
(2) Contradictory knowledge, wherein inconsistencies in CTI sources confuse models, causing unstable or conflicting predictions.
(3) Constrained generalization, wherein LLMs struggle to extend beyond familiar distributions, failing on emerging (zero-day) attack surfaces or novel TTPs.
@@
Our analysis further shows that these vulnerabilities directly undermine the effectiveness of techniques used in CTI pipelines. As summarized in Figure 2, each vulnerability spans multiple CTI stages and influences different aspects of LLM reasoning. For example, spurious correlations distort evidence retrieval, where LLMs amplify irrelevant co-occurrences in retrieved evidence, resulting in misguided contextualization and faulty attribution. Overall, the detailed studies reveal the root causes by combining vulnerabilities with the specific LLM workflows involved in different CTI stages, which provides an in-depth understanding of how these vulnerabilities propagate through the reasoning.
@@
By uncovering blind spots in LLM vulnerabilities for CTI, we provide foundational insights that can guide principled adaptations in future CTI-focused LLM systems. To facilitate follow-up research, we release codes at https://github.com/mengyuqiao/LLM-CTI.
@@ 2 Background and Motivation
2.1 Background: cyber threat intelligence stages
Motivating Example. A healthcare network detects suspicious outbound traffic linked to newly registered domains associated with QakBot. The security team first enriches event metadata by mapping proxy logs to known C2 infrastructure and retrieving prior reports for context. Next, they attribute QakBot’s reuse of infrastructure to the threat group TA577. Based on historical correlations, they then predict a likely transition to Cobalt Strike and eventual ransomware deployment. Finally, the team implements mitigation strategies, including generating Sigma detection rules and prioritizing patches for vulnerable Exchange servers.
As illustrated by the above motivating example, CTI practices are typically organized into a pipeline consisting of four stages, each of which involves distinct reasoning tasks and technical solutions:
Mitigation: Finally, CTI must support actionable decisions to reduce risk and guide incident response. This includes recommending specific patches, tuning detection signatures (e.g., YARA/Sigma rules (Alvarez, 2013; Roth & Patzke, 2017)), adjusting firewall or access control configurations, and drafting response playbooks. Involved techniques supporting this stage include mitigation mapping to associate observed TTPs or vulnerabilities with known defensive strategies, mitigation efficacy prediction to rank possible responses, and summarization to generate concise, structured remediation steps tailored to system environments.
Highlight . Involved techniques above not only define the operational CTI pipeline but also help explain the root causes of where vulnerabilities are likely to be introduced from threat landscape.
2.2 Motivation: LLMs remain insufficient in variant CTI tasks
@@
Benchmark CTI Scenario Coverage Unique Feature
CTIBench (Alam et al., 2024) General CTI tasks Multi-choice questions (MCQ)
SevenLLM-Bench (Ji et al., 2024) Report understanding Synthetic instances, MCQ, QA
SWE-Bench (Jimenez et al., 2023) Software bug fixing Program analysis & patching
CyberTeam (Liu et al., 2025) Blue-team threat hunting Open-ended decision-making
Evaluation Datasets. To evaluate LLM performance on CTI tasks, we leverage benchmarks (Alam et al., 2024; Ji et al., 2024; Jimenez et al., 2023; Liu et al., 2025) as well as real-world threat databases and platforms. The benchmarks are summarized in Table 1, which provide a broad coverage across all CTI stages and capture both structured (MCQ, QA) and unstructured (decision-making, patch generation) task formats. To unify their use, we standardize each instance into a CTI-oriented scenario explicitly aligned with one of the four CTI stages. For example, multi-choice questions (MCQs) from CTIBench and SevenLLM-Bench are reformulated into concrete threat hunting scenarios, such as analyzing a suspicious log entry to determine the relevant TTP or linking an IOC to a known malware family. Our preprocessing mitigates structural biases that could otherwise inflate LLM performance.
@@
CTI Task Metric Industry-leading, general-purpose Cybersecurity-specialized
@@
Evaluated LLMs. Our evaluation covers two complementary lines of models. First, we include industry-leading, general-purpose LLMs (e.g., GPT-5, Claude-Sonnet-4, Gemini-2.5), which represent state-of-the-art reasoning capabilities across domains. Second, we evaluate on open-source or API-accessible cybersecurity-specialized models (e.g., SecGPT (clouditera, 2025), DeepHat (DeepHat, 2025)), which are adapted to security operations through domain-specific training and curated CTI corpora. Investigating their performance gaps provides insights of vulnerabilities that cannot be fully addressed by either large-scale pretraining or cybersecurity-specialized adaptation. Appendix B.5 details all evaluated model and their versions.
@@
Model-specific Gap. Across extensive evaluations on a broad range of CTI tasks (detailed introductions in B.3), Table 2 shows model-specific gaps: general-purpose LLMs dominate understanding-heavy and synthesis tasks (e.g., populating attack graphs), while the best cyber-specialized agents may outstand on semantic-driven or operational outputs (e.g., patch recommendation). In contextualization, general models post sizable wins, e.g., Affected Systems F1 peaks at $\sim$0.82–0.88 versus $\sim$0.55–0.56 for most cyber agents; Source Reliability AUC tops out at $\sim$0.91 versus a best cyber score of $\sim$0.74, which reflects constrained long-context retrieval and instruction-following that also present in text generation tasks such as Threat Report Alignment and Event Timeline. In attribution and forecasting, the trend largely holds: general LLMs lead Threat Actor Linking (Acc up to $\sim$0.89) and Exploit Likelihood (AUC up to $\sim$0.86). Notably, targeted domain tuning can flip certain edges: a top cyber agent outperforms on Infrastructure Reuse (F1 $\sim$0.75 vs $\sim$0.68 best general), and cyber models consistently beat generals on operational ranking/classification such as Countermeasure Ranking (NDCG $\sim$0.62–0.63 vs $\sim$0.57–0.59), suggesting influences from environmental (enterprise) contexts. Performance among cyber agents is also more uneven (e.g., low F1 on False Flag Detection for some models), indicating sensitivity to the quality and coverage of training (or retrieving) evidence.
@@ 3 Methodology: Categorize Failure Instances
Overall, we do not fully trust LLM-as-judge or model-generated confidence to detect failure cases, due to the lack of transparency and the risk of “self-rationalized” reasoning. Nevertheless, the large scale of the evaluation set requires us to categorize instances efficiently beyond purely manual efforts. To address this, we propose a stratification approach that partitions instances based on their “failure” depth, followed by an autoregressive method combined with human-in-the-loop efforts to resolve the above RQs. Figure 3 illustrates the workflow. Algorithms are deferred to Appendix C.1
@@ 3.1 Stratifying instances without hard-label annotations (RQ1)
Most cyber threat intelligence tasks, such as mitigation rule generation, lack hard labels that allow a binary correct/incorrect assessment. However, reference materials are typically available from CTI reports and vulnerability advisories (e.g.,Cybersecurity and Infrastructure Security Agency (2024) (CISA); CVE Program (2024)). We therefore leverage the reference-based metrics (e.g., text similarity by BLEU) to quantify the matching degree between model outputs and authoritative sources.
Based on the calculated similarity scores, we stratify instances by their “failure” depth, ranging from severe mismatches to partial alignment and near matches. Specifically, we rank all instances and partition them into quantile-based strata (5% bins). Within each stratum, we initially inspect some “correct” and “failed” samples and record their score distributions into two groups. For the remaining instances, we classify them as “failed” if their scores fall within the range associated with failed samples; if they fall in the overlapping region between the two groups, we conduct additional manual inspection. We terminate the process once no new failure modes emerge (as defined in §3.2) and the distribution of failure modes (in §3.3) across strata converges to a stable ratio. Practically, this process is efficient as it requires manually inspecting no more than 3% of instances across all CTI tasks. For clarity, the detailed algorithm is provided in Algorithm 1.
3.2 Autoregressive failure mode determination (RQ2)
We avoid of using LLMs to directly determine failure modes (i.e., vulnerabilities), instead, we design an iterative process alternating between human annotation and LLM-assisted classification. Let $\mathcal{D}=\{x_{i}\}_{i=1}^{N}$ denote the set of failure instances (from stratification results in §3.1).

### v3

@@ Abstract
Large language models (LLMs) are increasingly used to help security analysts manage the surge of cyber threats, automating tasks from vulnerability assessment to incident response. Yet in operational CTI workflows, reliability gaps remain substantial. Existing explanations often point to generic model issues (e.g., hallucination), but we argue the dominant bottleneck is the threat landscape itself: CTI is heterogeneous, volatile, and fragmented. Under these conditions, evidence is intertwined, crowdsourced, and temporally unstable, which are properties that standard LLM-based studies rarely capture.
In this paper, we present a comprehensive empirical study of LLM vulnerabilities in CTI reasoning. We introduce a human-in-the-loop categorization framework that robustly labels failure modes across the CTI lifecycle, avoiding the brittleness of automated “LLM-as-a-judge” pipelines. We identify three domain-specific cognitive failures: spurious correlations from superficial metadata, contradictory knowledge from conflicting sources, and constrained generalization to emerging threats. We validate these mechanisms via causal interventions and show that targeted defenses reduce failure rates significantly. Together, these results offer a concrete roadmap for building resilient, domain-aware CTI agents.
@@ 1 Introduction
We are living in an era of rapid digital transformation, where technological advancements are tightly associated with the growing prevalence of cyber threats. In recent years, the cyber threat landscape has shifted dramatically, with reported Common Vulnerabilities and Exposures (CVEs) increasing by an average of 25% annually (ibm_cve_proliferation). In 2024 alone, more than 40,000 vulnerabilities were reported (cve-by-date). This surge can be attributed to the rising complexity of IT systems (qualys_cve_surge_2024), the widespread adoption of open-source software (dam2023towards; darkreading_cves_plan), and the accelerating pace of modern development cycles (ox_modernizing_appsec). Together, these dynamics expand the attack surface while making sole reliance on human analysts for vulnerability assessment and remediation increasingly infeasible.
@@
Large language models (LLMs) have recently demonstrated strong performance in a broad range of cyber threat intelligence (CTI) tasks. By adapting models through instruction fine-tuning or prompt-based automation, researchers have applied LLMs to support threat analysis and decision-making (zhang2023hackmentor), code vulnerability detection (du2024generalization), and defense against network intrusions (lavi2024fine). Despite these advances, substantial performance gaps are reported in their evaluations (deng2024pentestgpt; clairoux2024use; ji2024sevenllm; alam2024ctibench; liu2025benchmarking), suggesting that such limitations cannot be fully addressed through typical model adaptation or prompt-based automation. This raises a fundamental research question: What intrinsic vulnerabilities constrain the effectiveness of LLMs in supporting CTI tasks?
Figure 1: Failure ratios of LLM-based cybersecurity agents on large-scale CTI tasks (detailed evaluations in §2.2).
@@
Figure 2: Summarization and examples of vulnerability types of LLMs in various CTI stages. Ratios are calculated over the entire dataset (§2.2). Several vulnerabilities may present simultaneously in a single threat instance.
Although some vulnerabilities (e.g., hallucination (ji2023towards; martino2023knowledge; friel2023chainpoll), jailbreaking (niu2024jailbreaking; coalson2024prisonbreak)) arise from general deficiencies in LLM architecture and training methodologies that affect broad domains (aguilera2025llm), studying these limitations in a domain-agnostic manner fails to propose effective, cyber-centric solutions. Notably, developing LLM-based CTI agents requires navigating the threat landscape defined by the heterogeneity and volatility of its knowledge sources rigaki2024hackphyr; de2025cfa. Unlike static corpora, CTI is a fragmented ecosystem comprising structured databases (e.g., NVD (nvd)), semi-structured vendor advisories (berndtsson2017combining), and unstructured discussions on social media or dark web forums (wang2019automatic; rahman2020literature). Furthermore, cyber threats are ever-evolving; the severity or remediation status of a vulnerability is subject to rapid change based on real-time exploit availability (pureti2022zero). Consequently, standard LLM evaluations often fail to capture the complexity of the disparate, time-sensitive, and frequently contradictory threat landscape.
This work. Given that developing CTI agents requires reasoning under intertwined, crowdsourced, and imbalanced evidence, we aim to study: (1) which vulnerabilities are uniquely derived from the threat landscape itself; (2) how these factors introduce specific weaknesses in LLM-assisted CTI; and (3) what defensive solutions can effectively mitigate them. Specifically, we make the following contributions:
Example: CTI stages A healthcare network detects suspicious outbound traffic linked to newly registered domains associated with QakBot. The security team first enriches event metadata by mapping proxy logs to known C2 infrastructure and retrieving prior reports for context. Next, they attribute QakBot’s reuse of infrastructure to the threat group TA577. Based on historical correlations, they then predict a likely transition to Cobalt Strike and eventual ransomware deployment. Finally, the team implements mitigation strategies, including generating Sigma detection rules and prioritizing patches for vulnerable Exchange servers.
I. A comprehensive empirical evaluation. To characterize LLM failure modes in CTI, we conduct a large-scale, systematic study spanning the complete CTI lifecycle, from threat analysis to incident response. Our evaluation leverages a diverse suite of datasets, integrating established benchmarks (CTIBench (alam2024ctibench), SevenLLM-Bench (ji2024sevenllm), SWE-Bench (jimenez2023swe), and CyberTeam (liu2025benchmarking)) with real-world threat reports sourced from authoritative knowledge bases or threat feeds (e.g., (cve; nvd; CISA_KEV_Catalog)) to ensure comprehensiveness. Figure 1 shows summarized failure ratios among cyber LLM agents.
II. A failure categorization methodology. Based on our evaluation results, we develop a rigorous methodology to investigate and categorize failure instances. A critical challenge in analyzing extensive threat data is reliably scaling the classification of diverse failure modes. Standard LLM-as-a-judge approaches prove unreliable for automatic failure classification, as models tend to rationalize their own outputs or fail to critically evaluate contradictions (yamauchi2025empirical; guerdan2025validating). To address this, we propose an autoregressive, human-in-the-loop framework that categorizes failure instances with high efficiency and reliability (§3). This methodology enables us to extract data-grounded insights from large-scale evaluations while significantly reducing the manual annotation workload.
III. A taxonomy of CTI vulnerabilities and root cause analysis. Through large-scale experiments on industry-leading LLMs and LLM-powered cybersecurity agents, complemented by detailed case studies, we identify three dominant categories of vulnerabilities that limit LLM reasoning: (1) Spurious correlations, wherein LLMs over-attribute based on superficial or co-occurring features (e.g., mistaking commodity tools such as Mimikatz (Delpy_Mimikatz) as actor-specific evidence). (2) Contradictory knowledge, wherein inconsistencies in CTI sources confuse models, causing unstable or conflicting predictions. (3) Constrained generalization, wherein LLMs struggle to extend beyond familiar distributions, failing on emerging (zero-day) attack surfaces.
@@
Our analysis further shows that these vulnerabilities directly undermine the effectiveness of techniques used in CTI pipelines. As summarized in Figure 2, each vulnerability spans multiple CTI stages and influences different aspects of LLM reasoning. For example, spurious correlations distort evidence retrieval, where LLMs amplify irrelevant co-occurrences in retrieved evidence, resulting in misguided contextualization and faulty attribution. Similarly, contradictory knowledge weakens entity recognition and event extraction, as models produce inconsistent mappings of the same threat actor or IOC across conflicting sources, thereby breaking downstream attack chain construction. Overall, the detailed studies reveal the root causes by combining vulnerabilities with the specific LLM workflows involved in different CTI stages, which provides an in-depth understanding of how these vulnerabilities propagate through the reasoning.
IV. Validation and defensive strategies. We further validate the root causes of these vulnerabilities through a causal intervention study, showing that reasoning-time interventions can yield immediate performance gains. Building on these findings, we design and evaluate targeted mitigation strategies, including retrieval-based CTI knowledge filtering and inference-time reasoning constraints. Our results show that these lightweight, taxonomy-informed defenses consistently reduce failure rates, demonstrating that robust LLM-based CTI analysis requires protections explicitly tailored to the heterogeneity of threat landscape.
@@
In summary, this work uncovers blind spots in LLM vulnerabilities for CTI. We provide foundational insights and initial defensive efforts that can guide principled adaptations in future CTI-focused LLM systems. To facilitate follow-up research, we release codes at https://anonymous.4open.science/r/LLM-CTI-E773.
@@ 2 Background and Motivation
Benchmark Description Task Coverage Unique Feature
❶ Cont. ❷ Attr. ❸ Pred. ❹ Miti.
CTIBench (alam2024ctibench) General CTI tasks Multi-choice questions (MCQ)
SevenLLM-Bench (ji2024sevenllm) Report understanding Synthetic instances, MCQ, QA
SWE-Bench (jimenez2023swe) Software bug fixing Program analysis & patching
CyberTeam (liu2025benchmarking) Blue-team threat hunting Open-ended decision-making
@@
Detailed CTI Task Metric Industry-leading, general-purpose Cybersecurity-specialized
@@
2.1 Background: Stages of CTI Tasks
As illustrated by the example above, CTI practices are typically organized into a pipeline consisting of four stages, each of which involves distinct reasoning tasks and technical solutions:
Mitigation: Finally, CTI must support actionable decisions to reduce risk and guide incident response. This includes recommending specific patches, tuning detection signatures (e.g., YARA/Sigma rules (YARA; Sigma)), adjusting firewall or access control configurations, and drafting response playbooks. Involved techniques supporting this stage include mitigation mapping to associate observed TTPs or vulnerabilities with known defensive strategies, mitigation efficacy prediction to rank possible responses, and summarization to generate concise, structured remediation steps tailored to system environments.
Highlight . Involved techniques above not only define the operational CTI pipeline but also help explain the root causes of how vulnerabilities are triggered in cyberspace.
2.2 Motivation: LLMs remain insufficient in variant CTI tasks
Evaluation Datasets. To evaluate LLM performance on CTI tasks, we leverage benchmarks (alam2024ctibench; ji2024sevenllm; jimenez2023swe; liu2025benchmarking) as well as real-world threat databases and platforms. The benchmarks are summarized in Table 1, which provide a broad coverage across all CTI stages and capture both structured (MCQ, QA) and unstructured (decision-making, patch generation) task formats. To unify their use, we standardize each instance into a CTI-oriented scenario explicitly aligned with one of the four CTI stages. For example, multi-choice questions (MCQs) from CTIBench and SevenLLM-Bench are reformulated into concrete threat hunting scenarios, such as analyzing a suspicious log entry to determine the relevant TTP or linking an IOC to a known malware family. Our preprocessing mitigates structural biases that could otherwise inflate LLM performance.
@@ 2.2 Motivation: LLMs remain insufficient in variant CTI tasks
Evaluated LLMs. Our evaluation covers two complementary lines of models. First, we include industry-leading, general-purpose LLMs (e.g., GPT-5, Claude-Sonnet-4, Gemini-2.5), which represent state-of-the-art reasoning capabilities across domains. Second, we evaluate on open-source or API-accessible cybersecurity-specialized models (e.g., SecGPT (SecGPT), DeepHat (DeepHat)), which are adapted to security operations through domain-specific training and curated CTI corpora. Investigating their performance gaps provides insights of vulnerabilities that cannot be fully addressed by either large-scale pretraining or cybersecurity-specialized adaptation. Appendix B.5 details all evaluated model and their versions.
@@
Model-specific Gap. Across extensive evaluations on a broad range of CTI tasks (detailed introductions in B.3), Table 2 shows model-specific gaps: general-purpose LLMs dominate understanding-heavy and synthesis tasks (e.g., populating attack graphs), while the best cyber-specialized agents may stand out on semantic-driven or operational outputs (e.g., patch recommendation). In contextualization, general models post sizable wins, e.g., Affected Systems F1 peaks at $\sim$0.82–0.88 versus $\sim$0.55–0.56 for most cyber agents; Source Reliability AUC tops out at $\sim$0.91 versus a best cyber score of $\sim$0.74, which reflects constrained long-context retrieval and instruction-following that also present in text generation tasks such as Threat Report Alignment and Event Timeline. In attribution and forecasting, the trend largely holds: general LLMs lead Threat Actor Linking (Acc up to $\sim$0.89) and Exploit Likelihood (AUC up to $\sim$0.86). Notably, targeted domain tuning can flip certain edges: a top cyber agent outperforms on Infrastructure Reuse (F1 $\sim$0.75 vs $\sim$0.68 best general), and cyber models consistently outperform general LLMs on operational ranking/classification such as Countermeasure Ranking (NDCG $\sim$0.62–0.63 vs $\sim$0.57–0.59), suggesting influences from environmental (enterprise) contexts. Performance among cyber agents is also more uneven (e.g., low F1 on False Flag Detection for some models), implying sensitivity to the quality and coverage of training data.
@@ 3 Methodology: Categorize Failure Instances
Overall, we do not fully trust LLM-as-judge or model-generated confidence to detect failure cases, due to the lack of transparency and the risk of “self-rationalized” reasoning. Nevertheless, the large scale of the evaluation set requires us to categorize instances efficiently beyond purely manual efforts. To address this, we propose a stratification approach that partitions instances based on their “failure” depth, followed by an autoregressive method combined with human-in-the-loop efforts to resolve the above RQs. Figure 3 illustrates the overall workflow.
@@ 3.1 Stratifying instances without hard-label annotations (RQ1)
Input: Dataset $\mathcal{D}=\{x_{i}\}$, Models $\{y_{i}\}$, Refs $\{r_{i}\}$, Similarity measure $Sim(\cdot)$
Output: Failure Set $\mathcal{F}$
// 1. Score & Stratify
Compute scores $\mathcal{S}=\{s_{i}\mid s_{i}=Sim(y_{i},r_{i}),\forall i\}$
Sort $\mathcal{S}$ and partition into quantile strata $\mathcal{Q}=\{Q_{1},\dots,Q_{M}\}$ (bin size $\delta=0.05$)
// 2. Iterative Inspection
Initialize $\mathcal{F}\leftarrow\emptyset$, $Converged\leftarrow\text{False}$
foreach stratum $Q_{k}\in\mathcal{Q}$ while $\neg Converged$ do
  Sample anchors $A_{k}\subset Q_{k}$ for manual labeling
  Derive score ranges $R_{fail},R_{pass}$ from labeled anchors $A_{k}$
 // Propagate labels to remaining $x_{i}$
  foreach $x_{i}\in Q_{k}\setminus A_{k}$ do
     if $s_{i}\in R_{fail}$ then
        $\mathcal{F}\leftarrow\mathcal{F}\cup\{x_{i}\}$
     else if $s_{i}\in R_{pass}$ then
        continue
     else
        Manual check $x_{i}$; if failed $\mathcal{F}\leftarrow\mathcal{F}\cup\{x_{i}\}$
        // Boundary
     end if
  end foreach
 // 3. Termination Check
  Update failure mode distribution $D_{k}$
  if $D_{k}\approx D_{k-1}$ and no new modes found then
     $Converged\leftarrow\text{True}$
  end if
end foreach
return $\mathcal{F}$
Algorithm 1 Stratified Failure Sampling (RQ1)
Most cyber threat intelligence tasks, such as mitigation rule generation, lack binary hard labels. However, authoritative reference materials (e.g., advisories from CISA (CISA_KEV_Catalog) or CVE (cve)) allow us to leverage reference-based similarity metrics (e.g., BLEU) to quantify the alignment between model outputs and ground truth.
To efficiently identify failure cases, we design a stratification approach (Algorithm 1). We first compute similarity scores for all instances and partition them into 5% quantile-based strata. Within each stratum, we manually inspect a small set of anchor samples to establish score ranges for “correct” versus “failed” outputs. Remaining instances in the stratum are then automatically labeled based on these ranges; boundary cases falling in the overlapping region trigger additional manual review. This process proceeds iteratively through the strata and terminates once the distribution of failure modes stabilizes and no new modes emerge. Practically, this hybrid approach yields high-confidence failure sets while requiring manual inspection of less than 1.5% of the total instances.
Table 3: Distribution of categorized CTI vulnerabilities across different methods. Given the absence of ground truth (hard labels), human annotation serves as a near-gold standard.
Method Spurious Correlation Contradictory Knowledge Constrained Generalization
1.1 1.2 1.3 1.4 1.5 2.1 2.2 2.3 2.4 2.5 2.6 3.1 3.2 3.3 3.4
Co-men. Exploit. Confond. Skewed Hier. Temporal Conflict Semantic Diver. Misal. Counter. Distribut. Unseen Overfit. Env.
bias bias factor source metadata contra. report conflict data knowl. CTI bias pattern reasoning unaware
Human annotation 4.14 1.71 15.72 6.16 2.05 13.98 22.29 4.14 4.53 1.97 7.45 21.79 23.13 7.64 11.76
Clustering 8.53 5.24 10.16 12.41 6.83 18.37 15.68 10.21 9.87 6.54 14.26 18.59 16.42 15.84 16.23
Topic modeling 2.19 0.57 22.32 3.14 0.42 8.43 30.12 1.28 2.34 0.69 3.52 32.41 12.57 2.16 4.89
GPT judge 10.27 7.13 12.59 8.92 5.37 10.51 19.84 8.46 8.12 5.58 11.23 20.17 18.29 14.38 13.94
Gemini judge 5.82 3.42 14.23 7.56 3.29 15.19 21.07 5.53 5.86 3.13 9.47 21.54 26.83 11.27 11.62
Multi-agent debate 3.91 4.18 16.87 5.29 4.81 12.24 24.53 6.89 3.27 4.62 6.18 25.46 20.61 12.42 8.57
Ours 4.12 1.87 15.50 6.31 1.98 13.86 22.33 3.96 4.49 1.85 7.77 23.34 25.06 10.02 10.80
3.2 Failure model determination (RQ2)
We avoid using LLMs to directly determine failure modes (i.e., vulnerabilities), instead, we design an iterative process alternating between human annotation and LLM-assisted classification. Let $\mathcal{D}=\{x_{i}\}_{i=1}^{N}$ denote the set of failure instances (from stratification results in §3.1).